# Bayesian learning with Gaussian processes for low-dimensional representations of time-dependent nonlinear systems

### v1

@@ 2 Reduced-order models via operator inference
where $\mathbf{V}\in\mathbb{R}^{N\times r}$ is the basis matrix, $\hat{\mathbf{q}}(t)\in\mathbb{R}^{r}$ is the reduced state at time $t$, $\bar{\mathbf{q}}\in\mathbb{R}^{N}$ is, e.g., a steady-state solution or average state, and $r\ll N$ is the dimension of the reduced state. We typically require $\mathbf{V}$ to have orthonormal columns, in which case dynamics for $\hat{\mathbf{q}}(t)$ are derived by substituting eq. 2.2 into eq. 2.1 and multiplying both sides by $\mathbf{V}^{\mathsf{T}}$, resulting in the system
@@
In this work, we consider the scenario in which reduced-order operators such as $\hat{\mathbf{c}}$, $\hat{\mathbf{A}}$, $\hat{\mathbf{H}}$, and $\hat{\mathbf{B}}$ cannot be computed explicitly, as occurs when working with large, proprietary, or classified code bases, or in systems that must be transformed through a change of variables to first expose a polynomial structure similar to eq. 2.4a [49]. In these cases, the polynomial structure of $\mathbf{f}$ is known, but $\mathbf{f}$ itself is computationally inaccessible and cannot be used directly to form $\hat{\mathbf{f}}$.
@@
where $\hat{\mathbf{O}}\in\mathbb{R}^{r\times d(r,p)}$ is the operator matrix, $\mathbf{d}:\mathbb{R}^{r}\times\mathbb{R}^{p}\to\mathbb{R}^{d(r,p)}$ is the data vector, and $d(r,p)\in\mathbb{N}$ is the operator dimension. Although eq. 2.5 is posed in a linear form, the equations may be nonlinear in the reduced state $\hat{\mathbf{q}}(t)$ and/or the input $\mathbf{u}(t)$. For example, the quadratic system eq. 2.4b can be written as eq. 2.5 with $d(r,p)=1+r+r^{2}+p$ and
$\hat{\mathbf{O}}=[~{}\hat{\mathbf{c}}~{}~{}\hat{\mathbf{A}}~{}~{}\hat{\mathbf{H}}~{}~{}\hat{\mathbf{B}}~{}]\in\mathbb{R}^{r\times d(r,p)},\qquad\mathbf{d}(\hat{\mathbf{q}}(t),\hat{\mathbf{u}}(t))=\left[\begin{array}[]{c}1\\ \hat{\mathbf{q}}(t)\\ \hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)\\ \mathbf{u}(t)\end{array}\right]\in\mathbb{R}^{d(r,p)}.$ (2.6)
Given state-input pairs $\{(\mathbf{q}_{j},\mathbf{u}_{j})\}_{j=0}^{m-1}$ describing the solution of eq. 2.1 at $m$ discrete time instances $t_{0},\ldots,t_{m-1}$, OpInf learns an operator matrix $\hat{\mathbf{O}}$ by compressing the observed states as $\hat{\mathbf{q}}_{j}=\mathbf{V}^{\mathsf{T}}(\mathbf{q}_{j}-\bar{\mathbf{q}})\in\mathbb{R}^{r}$, then solving the regression
@@
Here, $\dot{\hat{\mathbf{q}}}_{j}\in\mathbb{R}^{r}$ is an estimate of the time derivative $\frac{\textrm{d}}{\textrm{d}t}\hat{\mathbf{q}}(t)=\mathbf{V}^{\mathsf{T}}\frac{\textrm{d}}{\textrm{d}t}{\mathbf{q}}(t)$ at time $t_{j}$, usually computed from finite differences of the reduced snapshots $\hat{\mathbf{q}}_{0},\ldots,\hat{\mathbf{q}}_{m-1}$, and the matrix $\mathbf{D}$ is called the data matrix.
@@ 3.1.1 Likelihood and prior
in which the mean function is assumed to be zero and $\kappa_{i}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is a kernel function specifying the covariance over the GP. If the kernel $\kappa_{i}$ is at least once differentiable in each argument, then the time-derivative of each $\hat{q}_{i}$ also follows a GP, and we have the joint processes
@@
Let ${\bf\it t}=(t_{0},\ldots,t_{m-1})^{\mathsf{T}}\in\mathbb{R}^{m}$ denote the $m\in\mathbb{N}$ time instances for which we have (reduced) snapshot data $\hat{\mathbf{q}}_{0},\ldots,\hat{\mathbf{q}}_{m-1}\in\mathbb{R}^{r}$. Define ${\bf\it Y}_{i}$ to be the random $m$-vector whose entries are the Gaussian random variables produced by evaluating the GP $\hat{q}_{i}$ at each time instance in ${\bf\it t}$, and let $\mathbf{y}_{i}\in\mathbb{R}^{m}$ denote the snapshot data for the $i$-th POD mode at times ${\bf\it t}$, that is,
@@
Note that $\mathbf{y}_{i}$ is an observation of ${\bf\it Y}_{i}$. Likewise, let ${\bf\it t}^{\prime}=(t_{0}^{\prime},\ldots,t_{m^{\prime}-1}^{\prime})^{\mathsf{T}}\in\mathbb{R}^{m^{\prime}}$ be the $m^{\prime}\in\mathbb{N}$ time instances (typically with $m^{\prime}\geq m$ and $[t_{0}^{\prime},t_{m^{\prime}-1}^{\prime}]\subset[t_{0},t_{m-1}]$) where the GP is to produce estimates, and define ${\bf\it Z}_{i}$ to be the random $m^{\prime}$-vector with entries given by the time derivative GP $\dot{\hat{q}}_{i}$ at times ${\bf\it t}^{\prime}$.
@@ 3.1.2 Posterior
Combining eq. 3.11eq. 3.12, Bayes’ rule gives the posterior distribution
@@ 3.1.3 Least squares formulation
Weight matrices $(\mathbf{W}_{1}^{zz})^{1/2},\ldots,(\mathbf{W}_{r}^{zz})^{1/2}\in\mathbb{R}^{d\times d}$,
@@ 3.2 Probabilistic reduced-order model predictions
For a given sample of $\hat{\mathbf{O}}$, we evaluate the reduced state $\hat{\mathbf{q}}(t)$ by solving the corresponding ROM eq. 2.5 over $[t_{0},t_{f}]$, which is computationally inexpensive due to its low dimensionality. The reduced state $\hat{\mathbf{q}}(t)$ therefore has the following distribution:
@@
in which all the $\hat{\mathbf{o}}_{i}$’s are marginalized. Here, the notation $p(\hat{\mathbf{q}}(t)|\,\hat{\mathbf{O}})$ indicates a deterministic model evaluation for a given instance of $\hat{\mathbf{O}}$; indeed, one can write the reduced state $\hat{\mathbf{q}}(t)$ as a function of $\hat{\mathbf{O}}$, i.e., $\hat{\mathbf{q}}(t;\hat{\mathbf{O}})$. Note, however, that the distribution of $\hat{\mathbf{q}}$ is not necessarily Gaussian, even though the posterior of $\hat{\mathbf{O}}$ is, primarily because $\hat{\mathbf{q}}(t;\hat{\mathbf{O}})$ is nonlinear with respect to $\hat{\mathbf{O}}$. Having constructed a posterior distribution for $\hat{\mathbf{O}}$, we use Monte Carlo sampling over the operator posterior to estimate the first- and second-order moments of $\hat{\mathbf{q}}(t)$, thereby establishing an uncertainty indicator for the model predictions.
@@ 4 Algorithmic procedure
The inference to define the posterior parameters eq. 3.16eq. 3.17 depends on the choice of kernel hyperparameters ${\bf\it\theta}_{1},\ldots,{\bf\it\theta}_{r}$, the data matrix $\tilde{\mathbf{D}}$, and the prior variance parameters ${\bf\it\gamma}_{1},\ldots,{\bf\it\gamma}_{r}$. This section provides strategies for selecting each of these quantities and details the associated numerical algorithms. For simplicity we return to the single-trajectory PDE setting, but the approach readily generalizes to cases with data from multiple trajectories (see Appendix B).
@@ 4.3 Prior variance selection
The vector ${\bf\it\gamma}_{i}\in\mathbb{R}^{d}$ parameterizing the prior distribution $\pi_{\texttt{prior}}(\hat{\mathbf{o}}_{i})$ plays the role of a Tikhonov regularizer in the generalized least-squares regression eq. 3.19 and strongly affects the stability of the resulting ROM. To methodically select ${\bf\it\gamma}_{1},\ldots,{\bf\it\gamma}_{r}$, we extend the approach introduced in [39], which balances model stability with accuracy in reproducing training data. For a candidate set $\{{\bf\it\gamma}_{i}\}_{i=1}^{r}$, we solve the corresponding Bayesian inference eq. 3.16eq. 3.17 for $\{({\bf\it\mu}_{i},{\bf\it\Sigma}_{i})\}_{i=1}^{r}$, sample $n_{s}\in\mathbb{N}$ times from the posterior operator distribution $\hat{\mathbf{O}}\sim[~{}\mathcal{N}({\bf\it\mu}_{1},{\bf\it\Sigma}_{1})~{}~{}\cdots~{}~{}\mathcal{N}({\bf\it\mu}_{r},{\bf\it\Sigma}_{r})~{}]^{\mathsf{T}}$, and solve the resulting reduced systems eq. 2.5 over the full time domain of interest $[t_{0},t_{f}]$. Recall that $t_{f}>t_{m-1}$, that is, the full simulation domain extends beyond where noisy data are observed. If the magnitude of any of the reduced model solutions exceeds a predetermined bound computed from the training data, the candidate set is disqualified for failing to produce a sufficiently stable model; otherwise, the sample mean of the model solutions is compared to the filtered state data $\tilde{\mathbf{y}}_{1},\ldots,\tilde{\mathbf{y}}_{r}$. A grid search, followed by a derivative-free optimization, is used to select optimal $\{{\bf\it\gamma}_{i}^{*}\}_{i=1}^{r}$ that minimizes the difference between the sample mean and the training data while not violating the stability criterion. Details are provided in Algorithm 2; in Section 5, the experiments use $n_{s}=20$ posterior samples and set the stability margin constant to $\varphi=5$. A similar strategy was deployed in our previous work [27]; however, in that work, only the MAP points $\hat{\mathbf{O}}=[~{}{\bf\it\mu}_{1}~{}~{}\cdots~{}~{}{\bf\it\mu}_{r}~{}]^{\mathsf{T}}$ were used to evaluate candidate regularizers. Our approach here, only slightly more computationally expensive, tests stability for multiple posterior samples and hence produces more robust models than the previous approach.

### v2

@@ 2 Reduced-order models via operator inference
where $\mathbf{V}\in\mathbb{R}^{N\times r}$ is the basis matrix, $\hat{\mathbf{q}}(t)\in\mathbb{R}^{r}$ is the reduced state at time $t$, $\bar{\mathbf{q}}\in\mathbb{R}^{N}$ is a fixed reference vector, and $r\ll N$ is the dimension of the reduced state. Both $\mathbf{V}$ and $\bar{\mathbf{q}}$, and hence the reduced state dimension $r$, are either inferred from training observations of the state $\mathbf{q}(t)$ or chosen based on problem-specific knowledge (e.g., $\bar{\mathbf{q}}$ may encode constant boundary conditions or represent a time-averaged solution). We typically require $\mathbf{V}$ to have orthonormal columns, in which case dynamics for $\hat{\mathbf{q}}(t)$ are derived by substituting eq. 2.2 into eq. 2.1 and multiplying both sides by $\mathbf{V}^{\mathsf{T}}$, resulting in the system
@@
In this work, we consider the scenario in which reduced-order operators such as $\hat{\mathbf{c}}$, $\hat{\mathbf{A}}$, $\hat{\mathbf{H}}$, and $\hat{\mathbf{B}}$ cannot be computed explicitly, as occurs when working with large, proprietary, or classified code bases, or in systems that must be transformed through a change of variables to first expose a polynomial structure similar to eq. 2.4a [52]. We will see examples of the latter in Section 5. In these cases, the polynomial structure of $\mathbf{f}$ is known, but $\mathbf{f}$ itself is computationally inaccessible and cannot be used directly to form $\hat{\mathbf{f}}$.
@@
where $\hat{\mathbf{O}}\in\mathbb{R}^{r\times d(r,p)}$ is the operator matrix, $\mathbf{d}:\mathbb{R}^{r}\times\mathbb{R}^{p}\to\mathbb{R}^{d(r,p)}$ is a known (typically nonlinear) function of the reduced state and input, and $d(r,p)\in\mathbb{N}$ is the operator dimension. The system eq. 2.5 is linear with respect to the operator matrix $\hat{\mathbf{O}}$ but is usually nonlinear in the reduced state $\hat{\mathbf{q}}(t)$ and/or the input $\mathbf{u}(t)$ due to $\mathbf{d}(\hat{\mathbf{q}}(t),\mathbf{u}(t))$. For example, the quadratic system eq. 2.4b can be written as eq. 2.5 with $d(r,p)=1+r+r^{2}+p$ and
$\hat{\mathbf{O}}=[~{}\hat{\mathbf{c}}~{}~{}\hat{\mathbf{A}}~{}~{}\hat{\mathbf{H}}~{}~{}\hat{\mathbf{B}}~{}]\in\mathbb{R}^{r\times d(r,p)},\qquad\mathbf{d}(\hat{\mathbf{q}}(t),\mathbf{u}(t))=\left[\begin{array}[]{c}1\\ \hat{\mathbf{q}}(t)\\ \hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)\\ \mathbf{u}(t)\end{array}\right]\in\mathbb{R}^{d(r,p)}.$ (2.6)
There are several ways to approach problems whose nonpolynomial nonlinear structure does not directly motivate the reduced structure eq. 2.5. Many problems can be posed in polynomial form after a change of variables [52], a strategy that has found application in several model reduction contexts (see, e.g., [39] for an example of variable transformation enabling balanced truncation ROMs). OpInf models have also been demonstrated to be effective in some scenarios where the true dynamics can only be partially represented through a polynomial form [21, 42, 51, 66]. Additionally, a large class of nonpolynomial nonlinear terms may be incorporated via discrete empirical interpolation [5].
Given state-input pairs $\{(\mathbf{q}_{j},\mathbf{u}_{j})\}_{j=1}^{m-1}$ describing the solution of (2.1) at $m$ discrete time instances $t_{0},\ldots,t_{m-1}$, a reference vector $\bar{\mathbf{q}}$, and a basis matrix $\mathbf{V}$, OpInf learns an operator matrix $\hat{\mathbf{O}}$ by compressing the observed states as $\hat{\mathbf{q}}_{j}=\mathbf{V}^{\mathsf{T}}(\mathbf{q}_{j}-\bar{\mathbf{q}})\in\mathbb{R}^{r}$, then solving the regression
@@
Here, $\dot{\hat{\mathbf{q}}}_{j}\in\mathbb{R}^{r}$ is an estimate of the time derivative $\frac{\textrm{d}}{\textrm{d}t}\hat{\mathbf{q}}(t)=\mathbf{V}^{\mathsf{T}}\frac{\textrm{d}}{\textrm{d}t}{\mathbf{q}}(t)$ at time $t_{j}$, usually computed from finite differences of the reduced snapshots $\hat{\mathbf{q}}_{0},\ldots,\hat{\mathbf{q}}_{m-1}$, and the matrix $\mathbf{D}$ is called the data matrix. Because $\mathbf{V}$, $\bar{\mathbf{q}}$, and the observations $\{(\mathbf{q}_{j},\mathbf{u}_{j})\}$ are known, the only unknowns in eq. 2.7 are the entries of the operator matrix $\hat{\mathbf{O}}$. Unless additional constraints on the unknown $\hat{\mathbf{O}}$ are imposed, eq. 2.7 is a linear least-squares problem since eq. 2.5 is linear with respect to $\hat{\mathbf{O}}$.
@@ 3.1.1 Likelihood and prior
in which the mean function is assumed to be zero and $\kappa_{i}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is a kernel function specifying the covariance over the GP. If the kernel $\kappa_{i}$ is at least once differentiable in each argument, then the time-derivative of each $\hat{q}_{i}$ also follows a GP, and we have the joint processes [55, 56]
@@
Let ${\bf\it t}=(t_{0},\ldots,t_{m-1})^{\mathsf{T}}\in\mathbb{R}^{m}$ denote the $m\in\mathbb{N}$ time instances for which we have (reduced) snapshot data $\hat{\mathbf{q}}_{0},\ldots,\hat{\mathbf{q}}_{m-1}\in\mathbb{R}^{r}$. Define ${\bf\it Y}_{i}=\hat{q}_{i}({\bf\it t})$ to be the random $m$-vector whose entries are the Gaussian random variables produced by evaluating the GP $\hat{q}_{i}$ at each time instance in ${\bf\it t}$, and let $\mathbf{y}_{i}\in\mathbb{R}^{m}$ denote the snapshot data for the $i$-th reduced state $\hat{q}_{i}$ at times ${\bf\it t}$, that is,
@@
Note that $\mathbf{y}_{i}$ is an observation of ${\bf\it Y}_{i}$. Likewise, let ${\bf\it t}^{\prime}=(t_{0}^{\prime},\ldots,t_{m^{\prime}-1}^{\prime})^{\mathsf{T}}\in\mathbb{R}^{m^{\prime}}$ be the $m^{\prime}\in\mathbb{N}$ time instances (typically with $m^{\prime}\geq m$ and $[t_{0}^{\prime},t_{m^{\prime}-1}^{\prime}]\subset[t_{0},t_{m-1}]$) where the GP is to produce estimates, and define ${\bf\it Z}_{i}=\dot{\hat{q}}_{i}({\bf\it t}^{\prime})$ to be the random $m^{\prime}$-vector with entries given by the time derivative GP $\dot{\hat{q}}_{i}$ at times ${\bf\it t}^{\prime}$.
@@
Remark 3.3 (Gaussianity preserved by affine reconstruction).
The independent reduced state GP representations $\hat{q}_{1},\ldots,\hat{q}_{r}$ of eq. 3.1 can be expressed together as a single vector-valued GP,
$\hat{\mathbf{q}}(t)\sim\text{vec-}\mathcal{GP}(\mathbf{0},\text{diag}(\kappa_{1}(t,t^{\prime}),\ldots,\kappa_{r}(t,t^{\prime}))\,.$ (3.12a)
If these reduced states are used in the approximation eq. 2.2, the reconstruction of full-order states is given by
$\mathbf{q}(t)\approx\mathbf{V}\hat{\mathbf{q}}(t)+\bar{\mathbf{q}}\sim\text{vec-}\mathcal{GP}(\bar{\mathbf{q}},\mathbf{V}\text{diag}(\kappa_{1}(t,t^{\prime}),\ldots,\kappa_{r}(t,t^{\prime}))\mathbf{V}^{\mathsf{T}})\,,$ (3.12b)
in which the covariance $\mathbf{V}\text{diag}(\kappa_{1}(t,t^{\prime}),\ldots,\kappa_{r}(t,t^{\prime}))\mathbf{V}^{\mathsf{T}}$ is a low-rank approximation of $\operatorname{Cov}[\mathbf{q}(t),\mathbf{q}(t^{\prime})]$. Hence, the GP-based modeling for the reduced states provides a quantification of uncertainties in the states on the full-order level as well.
@@ 3.1.2 Posterior
Combining eq. 3.11eq. 3.13, Bayes’ rule gives the posterior distribution
@@ 3.1.3 Least squares formulation
Weight matrices $(\mathbf{W}_{1}^{zz})^{1/2},\ldots,(\mathbf{W}_{r}^{zz})^{1/2}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$,
@@ 3.2 Probabilistic reduced-order model predictions
For a given sample of $\hat{\mathbf{O}}$, we evaluate the reduced state $\hat{\mathbf{q}}(t)$ by solving the corresponding ROM eq. 2.5. The reduced state $\hat{\mathbf{q}}(t)$ therefore has the following distribution:
@@
in which all the $\hat{\mathbf{o}}_{i}$’s are marginalized. Here, the notation $p(\hat{\mathbf{q}}(t)|\,\hat{\mathbf{O}})$ indicates a deterministic model evaluation for a given instance of $\hat{\mathbf{O}}$; indeed, one can write the reduced state $\hat{\mathbf{q}}(t)$ as a function of $\hat{\mathbf{O}}$, i.e., $\hat{\mathbf{q}}(t;\hat{\mathbf{O}})$. Note, however, that the distribution of $\hat{\mathbf{q}}$ is not necessarily Gaussian, even though the posterior of $\hat{\mathbf{O}}$ is, primarily because $\hat{\mathbf{q}}(t;\hat{\mathbf{O}})$ is nonlinear with respect to $\hat{\mathbf{O}}$. Having constructed a posterior distribution for $\hat{\mathbf{O}}$, we use Monte Carlo sampling over the operator posterior to estimate the moments of $\hat{\mathbf{q}}(t)$. Specifically, an individual sample $\hat{\mathbf{O}}_{k}$ defines an instance of the (deterministic) ROM eq. 2.5,
$\frac{\textup{d}}{\textup{d}t}\hat{\mathbf{q}}(t)=\hat{\mathbf{O}}_{k}\mathbf{d}(\hat{\mathbf{q}}(t),\mathbf{u}(t))\,,\quad\hat{\mathbf{q}}(t_{0})=\hat{\mathbf{q}}_{0}\,,\quad t\in[t_{0},t_{f}]\,,$
which is solved numerically for $\hat{\mathbf{q}}(t)$ over $[t_{0},t_{f}]$ using standard time-stepping integration methods. Each numerical solve is computationally inexpensive due to the low dimensionality of the system. After repeating this procedure for $k=1,\ldots,K$, the collection of $K$ trajectories is used to estimate the mean and higher-order moments of the reduced state at time $t$, thereby establishing an uncertainty indicator for the model predictions.
@@ 4 Algorithmic procedure
The inference to define the posterior parameters eq. 3.17eq. 3.18 depends on the choice of kernel hyperparameters ${\bf\it\theta}_{1},\ldots,{\bf\it\theta}_{r}$, the data matrix $\tilde{\mathbf{D}}$, and the prior variance parameters ${\bf\it\gamma}_{1},\ldots,{\bf\it\gamma}_{r}$. This section provides strategies for selecting each of these quantities and details the associated numerical algorithms. For simplicity we return to the single-trajectory PDE setting, but the approach readily generalizes to cases with data from multiple trajectories (see Appendix B).
@@ 4.3 Prior variance selection
The vector ${\bf\it\gamma}_{i}\in\mathbb{R}^{d}$ parameterizing the prior distribution $\pi_{\texttt{prior}}(\hat{\mathbf{o}}_{i})$ plays the role of a Tikhonov regularizer in the generalized least-squares regression eq. 3.20 and strongly affects the stability of the resulting ROM. To methodically select ${\bf\it\gamma}_{1},\ldots,{\bf\it\gamma}_{r}$, we extend the approach introduced in [42], which balances model stability with accuracy in reproducing training data. For a candidate set $\{{\bf\it\gamma}_{i}\}_{i=1}^{r}$, we solve the corresponding Bayesian inference eq. 3.17eq. 3.18 for $\{({\bf\it\mu}_{i},{\bf\it\Sigma}_{i})\}_{i=1}^{r}$, sample $n_{s}\in\mathbb{N}$ times from the posterior operator distribution $\hat{\mathbf{O}}\sim[~{}\mathcal{N}({\bf\it\mu}_{1},{\bf\it\Sigma}_{1})~{}~{}\cdots~{}~{}\mathcal{N}({\bf\it\mu}_{r},{\bf\it\Sigma}_{r})~{}]^{\mathsf{T}}$, and solve the resulting reduced systems eq. 2.5 over the full time domain of interest $[t_{0},t_{f}]$. Recall that $t_{f}>t_{m-1}$, that is, the full simulation domain extends beyond where noisy data are observed. If the magnitude of any of the reduced model solutions exceeds a predetermined bound computed from the training data, the candidate set is disqualified for failing to produce a sufficiently stable model; otherwise, the sample mean of the model solutions is compared to the filtered state data $\tilde{\mathbf{y}}_{1},\ldots,\tilde{\mathbf{y}}_{r}$. A grid search, followed by a derivative-free optimization, is used to select optimal $\{{\bf\it\gamma}_{i}^{*}\}_{i=1}^{r}$ that minimizes the difference between the sample mean and the training data while not violating the stability criterion. Details are provided in Algorithm 2; in Section 5, the experiments use $n_{s}=20$ posterior samples and set the stability margin constant to $\varphi=5$. A similar strategy was deployed in our previous work [29]; however, in that work, only the MAP points $\hat{\mathbf{O}}=[~{}{\bf\it\mu}_{1}~{}~{}\cdots~{}~{}{\bf\it\mu}_{r}~{}]^{\mathsf{T}}$ were used to evaluate candidate regularizers. Our approach, only slightly more computationally expensive, tests stability for multiple posterior samples and hence produces more robust models than the previous approach.